OpenStudy (anonymous):

Given that log_7(t) = 14.07, log_7(x) = -10.8, and log_7(z) = -17.02, find the following: log_7(tx^11)

OpenStudy (mathstudent55):

Remember these rules: \( \log (ab) = \log + \log b\) \(\log a^n = n \log a\) Now we apply those rules: \( \log_7 (tx^{11}) \) \(= \log_7 t + \log_7 x^{11} \) \(= \log_7 t + 11 \log_7 x\) Now just substitute the values of you have for the logs and evaluate.